OpenStudy (anonymous):

3. Another client, Wynona, decides that she will invest $5,000 per year in a 6% annuity for the first ten years, then $6,000 for the next ten years, and then $4,000 per year for the last ten years, how much will she accumulate? [Hint: Treat each ten-year period as as separate annuity and compute the Future Value. After the ten years, assume that the value will continue to grow at compound interest for the remaining years of the 30 years. Use tables from Unit 6 to compute compound interest

OpenStudy (anonymous):

5,000 * 0.06 *10 + 6,000 * 0.06 * 10 + 4,000 * 0.06 * 10 = 9000 thus $9,000

OpenStudy (sayan):

T=[5000(1+0.06)^10]+[6000(1)^10]+[4000(1)^10] =8954.238+6000+4000 =18954.238